# GrAviPaSt’s Lens to the Past: Unveiling the Evolution of Filamentary Structures

### v1

@@ Abstract
This paper examines the evolution of cosmic filaments across redshifts 1, 0.5, and 0 using the IllustrisTNG100-1 magneto-hydrodynamical simulation. To achieve this, we introduce GrAviPaSt, a simple and efficient filament identification method that leverages gravitational potential, an A*-like path-finding algorithm, and spanning trees. Applying this method to galaxy distributions at different redshifts allows us to analyze various filament properties, including their length, thickness, mass density contrast, and radial profile. Additionally, we investigate dynamic characteristics such as the mean distance of filament galaxies from the skeleton, their weighted mean velocity, and velocity trends normalized by their positions within the filaments. Our findings reveal the evolution of cosmic filaments from redshift 1 to 0, highlighting key differences across classifications. In particular, we examine the mass density contrast radial profile of filaments connecting two galaxy groups and those linking two galaxy clusters, identifying distinct differences in profile shape between these categories. Furthermore, in the context of weighted mean velocity, we analyze cosmic filaments exhibiting either negative or positive weighted mean velocity, demonstrating their differing evolutionary trends in terms of the mean distance of filament galaxies from the skeleton.
@@ 2 Data: Illustris-TNG
In this study, we selected all simulated galaxies from the illustrisTNG100-1 catalog that are brighter than approximately -16 in the r-band filter, possess stellar masses exceeding $10^{8}$ $M_{\odot}$, and exist at three different redshifts. We used the snapshots 50, 67 and 99, respectively corresponding to redshifts 1.0, 0.5 and 0. Using these thresholds, our final samples in these three redshifts contains about $34000$, $42000$ and $42000$ galaxies respectively.
@@ 3.1 Prim’s MST
GrAviPaSt method specifically uses Prim’s MST algorithm (R. C. Prim (1957)) instead of other MST algorithms such as Kruskal’s algorithm (e.g., as mentioned in I. Santiago-Bautista et al. (2020)). This choice is due to the optimal implementation of Prim’s MST. In GrAviPaSt method, the spatial coordinates of each galaxy group’s BCG (or BGG) serve as the node coordinates for the input of the MST algorithm. The Prim’s MST algorithm begins by initializing the process with an arbitrary node (vertex) from the input dataset, designating it as part of the growing minimum spanning tree (MST), while assigning an initial cost of infinity to all other nodes except the starting node, which has a cost of zero. Next, it iteratively selects the edge with the smallest weight that connects any node in the MST to a node outside it. For each newly added node, the algorithm updates the costs of its neighboring nodes, ensuring that any smaller edge weight replaces the current cost. This process is repeated until all nodes are included in the MST, thereby creating a structure with minimum-cost connections. Following the implementation of Prim’s MST, the pairs of galaxy groups between which the GrAviPaSt method will search for micro-filaments are identified. An example of the implementation of Prim’s MST is shown in Figure 1.
Figure 1: A schematic illustration depicting the sequential phases and steps of the GrAviPaSt method in identifying a macro-filament composed of three micro-filaments. Four galaxy groups are presented, with the method’s phases applied sequentially from bottom to top.
@@ 3.2 A*-like Path-Finder
Next, the GrAviPaSt method centers two spheres on the coordinates of the BCGs (or BGGs) of the two groups, using the virial radius $R_{200}$ of each group as the radius of the spheres. Subsequently, it tangentially places two pairs of planes around these spheres. Each pair of planes is parallel to each other but perpendicular to the other pair of planes. Using these four planes, the GrAviPaSt method constructs a pyramid-shaped structure around the two spheres, with the line connecting the two BCGs (or BGGs) forming the central axis. The algorithm retains only the region that is inside the pyramid, outside the spheres of the galaxy groups, and between the two spheres. The resulting space is a truncated pyramid: its base plane is tangential to the sphere of the larger group, its top plane is tangential to the sphere of the smaller group, and its four trapezoidal sides are defined by the two pairs of planes.
@@
$d=h\times(\frac{N}{V})^{\frac{1}{3}}$ (1)
where $h$ ($H_{0}/100$) is the Hubble constant, $N$ is the total number of galaxies in the input dataset (not the cropped box), and $V$ is the total volume of the input dataset. These grid points are shown between the galaxy group pairs in Figure 1.
@@
To identify the path with the deepest gravitational potential between two groups, the A*-like path-finding algorithm begins at the starting point. From this point, it considers all neighboring grid points that are closer to the BCG (or BGG) of the larger group than to the starting point. For each of these neighboring grid points, GrAviPaSt calculates the gravitational potential using Equation 2 and selects the grid point with the lowest (deepest) gravitational potential as the next step in the path. This process is repeated iteratively, with each subsequent point added to the path, until the algorithm reaches one of the ending point candidates. Once an ending point candidate is reached, the path-finding process halts.
@@ 4 Results
For the Prim’s MST phase of the GrAviPaSt method, the coordinates of BGGs (or BCGs) are chosen from galaxy groups containing at least 10 members. These selected coordinates serve as the nodes for the MST algorithm, with the micro-filaments identified in this study forming bridges between them.
@@ 4.1 Micro-Filaments
In this subsection, we examine the characteristics of the micro-filaments identified using the GrAviPaSt method. First, we analyze the length and thickness distributions of these filaments, providing insights into their structural properties. Subsequently, we investigate the respected velocity of filament galaxies along the micro-filament skeleton, revealing their dynamic behavior. Finally, we explore the radial profiles of the micro-filaments to identify patterns in their density and structural composition.
@@ 4.1.1 Length Evolution
The statistical analysis, performed using two-sided Kolmogorov–Smirnov (K-S) tests, further supports these findings. Comparisons of the co-moving length distributions of micro-filaments at redshifts 0 and 0.5, 0 and 1, as well as 0.5 and 1, yield p-values of 0.92, 0.68, and 0.52, respectively. These values indicate no statistically significant differences among the distributions. In contrast, comparisons of the proper length distributions result in p-values of $1.8\times 10^{-5}$, $9.34\times 10^{-15}$, and $1.21\times 10^{-4}$ respectively, indicating statistically significant differences among these distributions. Additionally, Figure 2 illustrates that the micro-filament population, in both proper and co-moving coordinates, is consistently dominated by shorter filaments across all redshifts.
@@ 4.1.2 Thickness Evolution
Further statistical analysis using K-S tests reinforces these findings. Comparisons of co-moving thickness distributions at redshifts 0 and 0.5, 0 and 1, and 0.5 and 1 yield p-values of 0.99, 0.23, and 0.22, respectively, indicating no statistically significant differences among the distributions, consistent with the observations for micro-filament length distributions. Conversely, comparisons in proper coordinates result in p-values of $8.15\times 10^{-10}$, $1.88\times 10^{-39}$, and $3.57\times 10^{-13}$, demonstrating statistically differences among these distributions.
@@ 4.1.3 Filament Galaxies Respected Velocity
It is important to note that the micro-filament skeletons, as identified by the GrAviPaSt method, begin at the smaller galaxy group and extend to the larger one (Section 3.2). Smaller normalized positions correspond to galaxies near the smaller group, while larger normalized positions indicate galaxies closer to the larger group. In the bottom panel of Figure 2, negative velocities represent galaxies moving toward the micro-filament skeleton, while positive velocities indicate galaxies moving away from it. As shown in this figure, galaxies with smaller normalized positions, and therefore closer to the smaller group, tend to exhibit more negative velocities and approaching the filament skeleton. Moving further along the micro-filament skeleton, galaxy velocities gradually become more positive. Within the normalized range of 0 to 0.3, the median velocities are negative across different redshifts, transitioning to values close to zero around 0.3. Beyond this point, within the range of 0.3 to 1, the median velocities remain approximately zero.
@@ 4.1.4 Radial Profiles
Figure 3: The radial profiles of mass density contrast are presented for total (top), CC (middle) and GG micro-filaments (bottom), each displayed across three redshifts using different colors. The corresponding fit curves, based on the Plummer profile, are represented by dashed lines.
Beyond the total population of micro-filaments, we also examined two specific types: those connecting galaxy clusters (CCs) and those linking groups of galaxies (GGs). CC micro-filaments were identified as structures bridging galaxy groups with at least 100 members, while GG micro-filaments connected groups containing 10 to 20 members. The middle and bottom panels of Figure 3 illustrate the profiles and corresponding fit-curves for CC and GG micro-filaments, modeled using the Plummer profile. Their fitted parameters and MSEs are detailed in Table 1.
As shown in Figure 3, the radial mass density profiles of CC micro-filaments across all three redshifts exhibit a general shape similar to that of the overall micro-filament population. In contrast, the profiles of GG micro-filaments differ significantly from those of both total filaments and CC micro-filaments.
@@
Parameter $\alpha$ $\beta$ $\gamma$ $MSE$ $\alpha$ $\beta$ $\gamma$ $MSE$ $\alpha$ $\beta$ $\gamma$ $MSE$
Total 0.79 4.81 1.41 0.0005 0.82 4.49 1.32 0.0002 0.56 6.97 2.07 0.0005
CCs 0.87 5.56 1.70 0.0015 0.47 6.91 2.38 0.0007 0.77 7.27 2.27 0.0044
GGs 0.37 5.79 1.69 0.0098 0.14 6.89 2.25 0.0019 0.16 7.38 2.49 0.0053
@@ 4.2.1 Effective Length Evolution
As observed, the range of length distributions differs across redshifts, with macro-filaments gradually increasing in length as redshift decreases. In all three redshifts, shorter macro-filaments make up the majority of the population; however, their dominance diminishes as redshift decreases. Kernel Density Estimation (KDE) curves are overlaid on the length histograms, showing peaks at 6.86, 4.59, and 3.98 for redshifts 0, 0.5, and 1, respectively. These results indicate evolution in the length distributions over cosmic time. Additionally, a two-sided K-S test conducted on the length distributions of macro-filaments between redshifts 0 and 0.5, 0 and 1, 0.5 and 1, yields p-values of 0.06, $4.85\times 10^{-6}$, and $1.6\times 10^{-3}$ respectively, indicating a statistically significant difference between the length distributions at redshifts 0 and 0.5 compared to that at redshift 1.
Figure 4: The distribution of macro-filaments length in proper coordinates and mass density contrast across three redshifts is shown. Scatter plots and the mass density contrast histograms use distinct colors to represent the two macro-filament populations, categorized by the number of filament galaxies they host. Horizontal lines mark the median length within each population’s distribution, while vertical lines represent the median mass density contrast at each redshift.
@@ 4.2.2 Mass Density Contrast Evolution
As illustrated in Figure 4, the mass density contrast distribution remains nearly consistent across different redshifts for both groups. The median of each group exhibits a gradual decline as redshift decreases, primarily due to the thickening of micro-filaments, as discussed earlier (Section 4.1.2). Additionally, the gap between the medians of the two groups widens with decreasing redshift. A statistical analysis using a two-sided K-S test on the mass density contrast distributions of macro-filaments with a lower galaxy population ($\leq 10$) produced p-values of 0.74, 0.02, and 0.25 for comparisons between redshifts 0 and 0.5, 0 and 1, and 0.5 and 1, respectively. These results suggest no significant difference between the distributions at redshifts 0 and 0.5, as well as between 0.5 and 1; however, a statistically significant difference is observed between redshifts 0 and 1. For macro-filaments with a higher galaxy population ($>10$), the same K-S tests yielded p-values of 0.98, 0.79, and 0.76 for the same redshift comparisons. These findings indicate no significant differences in mass density contrast distributions of this group across the three redshifts.
Table 2: Total luminosity ($log_{10}$ $L_{\odot}$), mass ($log_{10}$ $M_{\odot}$) and volume of the macro-filaments in both co-moving ($(Mpc/h)^{3}$) and proper ($Mpc^{3}$) through different redshifts.
@@
Total mass $1.74\times 10^{4}$ $1.83\times 10^{4}$ $1.54\times 10^{4}$
Total luminosity $1.75\times 10^{4}$ $1.85\times 10^{4}$ $1.56\times 10^{4}$
Total proper volume 1415.82 603.53 242.05
Total co-moving volume 440.09 428.21 423.11

### v2

@@ Abstract
This paper examines the evolution of cosmic filaments across redshifts 1, 0.5, and 0 using the IllustrisTNG100-1 magneto-hydrodynamical simulation. To achieve this, we introduce GrAviPaSt, a simple, efficient and parameter-free filament identification method that leverages gravitational potential, an A*-like path-finding algorithm, and spanning trees. Applying this method to galaxy distributions at different redshifts allows us to analyze various filament properties, including their length, thickness, mass density contrast, and radial profile. Additionally, we investigate dynamic characteristics such as the mean distance of filament galaxies from the skeleton, their weighted mean velocity, and velocity trends normalized by their positions within the filaments. Our findings reveal the evolution of cosmic filaments from redshift 1 to 0, highlighting key differences across classifications. In particular, we examine the mass density contrast radial profile of filaments connecting two galaxy groups and those linking two galaxy clusters, identifying distinct differences in profile shape between these categories. Furthermore, in the context of weighted mean velocity, we analyze cosmic filaments exhibiting either negative or positive weighted mean velocity, demonstrating their differing evolutionary trends in terms of the mean distance of filament galaxies from the skeleton.
@@ 2 Data: Illustris-TNG
In this study, we selected all simulated galaxies from the illustrisTNG100-1 catalog that are brighter than approximately -16 in the r-band filter, possess stellar masses exceeding $10^{8}$ $M_{\odot}$, and exist at three different redshifts. We used the snapshots 50, 67 and 99, respectively corresponding to redshifts 1.0, 0.5 and 0. Using these thresholds, our final samples in these three redshifts contains about $46000$, 4500045000 and $41000$ galaxies respectively.
@@ 3.1 Prim’s MST
GrAviPaSt method specifically uses Prim’s MST algorithm (R. C. Prim (1957)) instead of other MST algorithms such as Kruskal’s algorithm (e.g., as mentioned in I. Santiago-Bautista et al. (2020)). This choice is due to the optimal implementation of Prim’s MST. In GrAviPaSt method, the spatial coordinates of each galaxy group’s BGG (or BCG) serve as the node coordinates for the input of the MST algorithm. The Prim’s MST algorithm begins by initializing the process with an arbitrary node (vertex) from the input dataset, designating it as part of the growing minimum spanning tree (MST), while assigning an initial cost of infinity to all other nodes except the starting node, which has a cost of zero. Next, it iteratively selects the edge with the smallest weight that connects any node in the MST to a node outside it. For each newly added node, the algorithm updates the costs of its neighboring nodes, ensuring that any smaller edge weight replaces the current cost. This process is repeated until all nodes are included in the MST, thereby creating a structure with minimum-cost connections. Following the implementation of Prim’s MST, the pairs of galaxy groups between which the GrAviPaSt method will search for micro-filaments are identified. An example of the implementation of Prim’s MST is shown in Figure 1.
Figure 1: A schematic illustration depicting the sequential phases and steps of the GrAviPaSt method in identifying a macro-filament composed of three micro-filaments. Four galaxy groups are presented, with the method’s phases applied sequentially from bottom to top. The black lines connecting the centers of the galaxy groups illustrate the application of Prim’s MST algorithm, used to identify the galaxy group pairs between which GrAviPaSt searches for micro-filaments. Between the first and second galaxy groups, the schematic depicts the A*-like path-finder algorithm that identifies the path of deepest gravitational potential using grid points. Between the second and third, as well as the third and fourth galaxy groups, filament galaxies and the micro-filament radius are identified. Finally, the angles $\theta$ between the connection points of the micro-filaments attached to each galaxy group are illustrated. These angles are used to chain together micro-filaments in order to construct the macro-filament.
@@ 3.2 A*-like Path-Finder
Next, the GrAviPaSt method centers two spheres on the coordinates of the BGGs (or BCGs) of the two groups, using the virial radius $R_{200}$ of each group as the radius of the spheres. Subsequently, it tangentially places two pairs of planes around these spheres. Each pair of planes is parallel to each other but perpendicular to the other pair of planes. Using these four planes, the GrAviPaSt method constructs a pyramid-shaped structure around the two spheres, with the line connecting the two BGGs (or BCGs) forming the central axis. The algorithm retains only the region that is inside the pyramid, outside the spheres of the galaxy groups, and between the two spheres. The resulting space is a truncated pyramid: its base plane is tangential to the sphere of the larger group, its top plane is tangential to the sphere of the smaller group, and its four trapezoidal sides are defined by the two pairs of planes.
@@
$d=\sqrt[3]{\frac{N}{V}}$ (1)
where $N$ is the total number of galaxies in the input dataset (not the cropped box), and $V$ is the total volume of the input dataset. These grid points are shown between the galaxy group pairs in Figure 1.
@@
To identify the path with the deepest gravitational potential between two groups, the A*-like path-finding algorithm begins at the starting point. From this point, it considers all neighboring grid points that are closer to the BGG (or BCG) of the larger group than to the starting point. For each of these neighboring grid points, GrAviPaSt calculates the gravitational potential using Equation 2 and selects the grid point with the lowest (deepest) gravitational potential as the next step in the path. This process is repeated iteratively, with each subsequent point added to the path, until the algorithm reaches one of the ending point candidates. Once an ending point candidate is reached, the path-finding process halts.
@@ 4 Results
For the Prim’s MST phase of the GrAviPaSt method, the coordinates of BGGs (or BCGs) are chosen from galaxy groups containing at least 5 members. These selected coordinates serve as the nodes for the MST algorithm, with the micro-filaments identified in this study forming bridges between them.
@@ 4.1 Micro-Filaments
In this subsection, we examine the characteristics of the micro-filaments. First, we analyze the length and thickness distributions of these filaments, providing insights into their structural properties. Subsequently, we investigate the respected velocity of filament galaxies along the micro-filament skeleton, revealing their dynamic behavior. Finally, we explore the radial profiles of the micro-filaments to identify patterns in their density and structural composition.
@@ 4.1.1 Length Evolution
Table 1: P-values from two-sided Kolmogorov–Smirnov tests comparing the distributions of micro-filament characteristics between pairs of redshifts.
Redshift pairs 0 and 0.5 0 and 1 0.5 and 1
Length Co-moving 0.32 0.02 0.16
Proper $1.46\times 10^{-7}$ $5.13\times 10^{-29}$ $3.15\times 10^{-11}$
Thickness Co-moving $9.03\times 10^{-3}$ 0.41 0.03
Proper $1.18\times 10^{-4}$ $3.44\times 10^{-36}$ $2.19\times 10^{-25}$
The statistical analysis, performed using two-sided Kolmogorov–Smirnov (K–S) tests, further supports these findings. The resulting p-values, summarized in Table 1, correspond to comparisons of the co-moving length distributions of micro-filaments at redshifts 0 and 0.5, 0 and 1, and 0.5 and 1. These results indicate no statistically significant differences between the distributions at redshifts 0 and 0.5, as well as 0.5 and 1, but show a significant difference between redshifts 0 and 1. In contrast, comparisons of the proper length distributions, also yield p-values listed in Table 1, which indicate statistically significant differences among these distributions. Additionally, Figure 2 illustrates that the micro-filament population, in both proper and co-moving coordinates, is consistently dominated by shorter filaments across all redshifts.
@@ 4.1.2 Thickness Evolution
Further statistical analysis using K–S tests reinforces these findings. The resulting p-values, summarized in Table 1, correspond to comparisons of the co-moving thickness distributions at redshifts 0 and 0.5, 0 and 1, and 0.5 and 1. These values indicate statistically significant differences between the distributions at redshifts 0 and 0.5, as well as 0.5 and 1, but no significant difference between redshifts 0 and 1, in contrast to the observations for micro-filament length distributions (4.1.1). Comparisons in proper coordinates also yield p-values listed in Table 1, demonstrating statistically differences among these distributions.
@@ 4.1.3 Filament Galaxies Respected Velocity
It is important to note that the micro-filament skeletons, as identified by the GrAviPaSt method, begin at the smaller galaxy group and extend to the larger one (Section 3.2). Smaller normalized positions correspond to galaxies near the smaller group, while larger normalized positions indicate galaxies closer to the larger group. In the bottom panel of Figure 2, negative velocities represent galaxies moving toward the micro-filament skeleton, while positive velocities indicate galaxies moving away from it. As shown in this figure, the median of the respected velocity of filament galaxies remains close to zero across different redshifts. The trend is nearly identical at redshifts 1 and 0.5, but a subtle shift toward more negative velocities is observed from redshift 0.5 to 0.
@@ 4.1.4 Radial Profiles
Figure 3: The radial profiles of mass density contrast are presented for total (top), $C$ (middle) and GG micro-filaments (bottom), each displayed across three redshifts using different colors. The corresponding fit curves, based on the Plummer profile, are represented by dashed lines.
Beyond the total population of micro-filaments, we also examined two specific types: those connecting galaxy clusters (CCs) and those linking groups of galaxies (GGs). CC micro-filaments were identified as structures bridging galaxy groups with at least 50 members, while GG micro-filaments connected groups containing 5 to 10 members. The middle and bottom panels of Figure 3 illustrate the profiles and corresponding fit-curves for CC and GG micro-filaments, modeled using the Plummer profile. Their fitted parameters and MSEs are detailed in Table 2.
As shown in Figure 3, the radial mass density profiles of CC and GG micro-filaments across all three redshifts differ significantly from those of total filaments, indicating distinct evolutionary trajectories for these micro-filament types from redshift 1 to 0.
@@
Parameter $\alpha$ $\beta$ $\gamma$ MSEMSE $\alpha$ $\beta$ $\gamma$ MSEMSE $\alpha$ $\beta$ $\gamma$ MSEMSE
Total 0.73 4.95 1.44 0.0006 0.68 5.55 1.62 0.0007 0.84 4.71 1.36 0.0008
$C$s 0.63 7.15 2.22 0.0016 0.61 7.08 2.31 0.0173 0.71 4.68 1.88 0.0739
$G$s 0.29 5.31 1.57 0.0096 0.34 5.51 1.63 0.0008 0.00015 10.01 6.13 0.0085
@@ 4.2.1 Effective Length Evolution
As observed, the range of length distributions differs across redshifts, with macro-filaments gradually increasing in length as redshift decreases. In all three redshifts, shorter macro-filaments make up the majority of the population; however, their dominance diminishes as redshift decreases. Kernel Density Estimation (KDE) curves are overlaid on the length histograms, showing peaks at 5.59, 4.54, and 2.87 for redshifts 0, 0.5, and 1, respectively. These results indicate evolution in the length distributions over cosmic time. Additionally, a two-sided K-S test conducted on the length distributions of macro-filaments between redshifts 0 and 0.5, 0 and 1, and 0.5 and 1. The resulting p-values, summarized in Table 3, indicate statistically significant differences among these distributions.
Figure 4: The distribution of macro-filaments length in proper coordinates and mass density contrast across three redshifts is shown. Scatter plots and the mass density contrast histograms use distinct colors to represent the two macro-filament populations, categorized by the number of filament galaxies they host. Horizontal lines mark the median length within each population’s distribution, while vertical lines represent the median mass density contrast.
Table 3: P-values from two-sided Kolmogorov–Smirnov tests comparing the distributions of macro-filament characteristics between redshift pairs. Here, NN and $\bar{V}$ are the population count and the weighted mean velocity of filament galaxies, respectively.
Redshift pairs 0 and 0.5 0 and 1 0.5 and 1
Effective length $2.21\times 10^{-3}$ $1.47\times 10^{-11}$ $1.83\times 10^{-5}$
Mass density contrast $N\leq 10$ $1.51\times 10^{-17}$ $4.02\times 10^{-40}$ $3.29\times 10^{-13}$
$N>10$ $6.93\times 10^{-11}$ $2.11\times 10^{-19}$ $2.86\times 10^{-7}$
Mean distance to skeleton $\bar{V}\leq 0$ $2.11\times 10^{-3}$ $8.84\times 10^{-19}$ $4.39\times 10^{-14}$
$\bar{V}>0$ $5.52\times 10^{-5}$ $1.13\times 10^{-19}$ $2.71\times 10^{-16}$
@@ 4.2.2 Mass Density Contrast Evolution
As shown in Figure 4, the mass density contrast distribution is more concentrated at redshift 1 and becomes increasingly dispersed as redshift decreases. The median value of each group consistently rises with decreasing redshift. Furthermore, the difference between the medians of the two groups, which is evident at redshift 1, disappears at redshift 0.5, and the medians remain close at redshifts 0.5 and 0. This trend suggests that the poor population group undergoes a greater increase in mass density contrast compared to the rich population group.
A statistical analysis using two-sided K–S tests was conducted on the mass density contrast distributions of macro-filaments with a lower galaxy population ($\leq 10$). The resulting p-values, presented in Table 3, correspond to comparisons between redshifts 0 and 0.5, 0 and 1, and 0.5 and 1, and indicate statistically significant difference between distributions at different redshifts. For macro-filaments containing a larger galaxy population (>10>10), the same K–S tests yield p-values also listed in Table 3, which similarly confirm significant differences in the mass density contrast distributions across all three redshifts.
Table 4: Total luminosity ($log_{10}$ $L_{\odot}$), mass ($log_{10}$ $M_{\odot}$) and volume of the macro-filaments in both co-moving ($(Mpc/h)^{3}$) and proper ($Mpc^{3}$) coordinates through different redshifts.
@@
Total mass $1.51\times 10^{4}$ $1.96\times 10^{4}$ $1.88\times 10^{4}$
Total luminosity $1.53\times 10^{4}$ $2.05\times 10^{4}$ $2.00\times 10^{4}$
Total proper volume 950.29 658.44 227.63
Total co-moving volume 295.39 467.16 397.89